# Boosting Active Defense Persistence: A Two-Stage Defense Framework Combining Interruption and Poisoning Against Deepfake

### v2

@@ IV-D Analysis of Defense Persistence
This section is dedicated to the core contribution of this work, which is to validate the persistence of our TSDF framework. We conduct a direct comparative analysis that first establishes the failure of conventional SOTA defenses against an adaptive attack. We then demonstrate that our framework not only withstands the same attack but also enhances its defensive capabilities, thereby proving its superior persistence.
@@
The results presented in Table III reveal a critical vulnerability in conventional interruption-only methods. For instance, with the CMUA method, the average SSIM score on the CelebA dataset increases from 0.64 to 0.79 after retraining, while the average FID score falls from 209.46 to 127.42. The rising SSIM indicates that the output of perturbed examples have become more similar to the original images, while the falling FID confirms they have become more realistic and less distorted. Both metrics prove that the interruption effect has been largely nullified. This degradation provides clear quantitative evidence that adaptive attacks severely compromise conventional interruption-only defenses, proving their lack of persistence. This highlights the key vulnerability our TSDF framework is designed to solve.
@@
To provide a direct comparison against this established baseline of failure, we subjected our TSDF framework to the identical adaptive attack.
To validate the persistence of our framework, we conduct a direct comparative experiment against the SOTA interruption methods. To ensure a fair comparison, we evaluate all methods on the same two representative models, StarGAN and AttGAN. We simulated an adaptive attack by retraining these models on datasets protected by each respective defense. For TSDF, the retraining dataset was first processed by the poisoning mechanism. We then measured the interruption effectiveness of each defense before and after this adaptive attack to quantify their persistence.
The results presented in Table IV provide clear quantitative evidence of TSDF’s superior persistence. After retraining, the defenses of both CMUA and FOUND degraded significantly. CMUA’s SSIM score increased from 0.59 to 0.77, and its FID score dropped from 243.00 to 122.46. These results confirm that the retrained models successfully learned to bypass these perturbations. In contrast, the output from the model retrained on TSDF-protected data become even more distorted. Its SSIM score decreases from 0.46 to 0.31, and its FID score increases from 283.01 to 348.65. This outcome is not because the interruption perturbation became stronger. It is direct evidence of the poisoning component’s success in degrading the quality of deepfake models.
This enhanced persistence is attributed to the efficacy of the poisoning component. As detailed in the model discrepancy analysis in Table V, retraining on TSDF data fundamentally destroys the deepfake model’s own generative capabilities. For instance, a StarGAN model’s FID score on clean inputs after this process increases from a baseline of 43.77 to 152.93. Since the model’s core ability to generate quality images is compromised, it cannot learn to properly adapt to or bypass the complex interruption perturbation.
@@ IV-E Visual and Qualitative Analysis
A crucial aspect of our framework is its ability to provide strong defense while remaining invisible to the human eye. Fig. 6 provides a direct visual comparison of an original image against images protected by CMUA, FOUND and TSDF. The visual differences are nearly imperceptible, demonstrating the excellent concealment of our perturbation. This qualitative assessment is supported by the quantitative data in Table VI, which shows high PSNR and SSIM scores for the protected images across all datasets. For instance, on the CelebA dataset, TSDF achieves a PSNR of 33.57 and an SSIM of 0.90, demonstrating that our framework can provide a defense without compromising image imperceptibility.
@@ IV-F Ablation Study
We perform ablation studies to analyze the contribution of key components of the TSDF framework. Specifically, we evaluate the effectiveness of the intensity separation mechanism and the Feature Enhancement strategy.
TABLE VII: Impact of Perturbation Threshold on Defense Performance. This table evaluates how different threshold settings affect the trade-off between the two defensive functions. The results show the performance across various deepfake models and face detectors. Blank entries indicate that the corresponding defensive effect was non-existent at that threshold.
@@
0.0 L2mask↑ 0.18 0.13 0.65 0.29 - - -
SRmask↑ 98.98% 96.06% 100.00% 100.00% - - -
PSNR↓ 16.75 17.67 7.17 13.90 - - -
SSIM↓ 0.78 0.64 0.24 0.57 - - -
@@
0.2 L2mask↑ 0.18 0.13 0.62 0.26 - - -
SRmask↑ 98.72% 94.56% 100.00% 100.00% - - -
PSNR↓ 17.12 18.20 7.31 13.98 - - -
SSIM↓ 0.78 0.66 0.25 0.57 - - -
F1-score↓ - - - - 0.81 0.87 0.78
0.3 L2mask↑ 0.17 0.12 0.63 0.29 - - -
SRmask↑ 98.25% 92.55% 100.00% 100.00% - - -
PSNR↓ 17.66 18.68 7.56 13.97 - - -
SSIM↓ 0.79 0.68 0.24 0.57 - - -
F1-score↓ - - - - 0.78 0.78 0.76
0.4 L2mask↑ 0.20 0.02 0.70 0.29 - - -
SRmask↑ 98.90% 0.06% 100.00% 100.00% - - -
PSNR↓ 16.76 28.13 7.21 13.82 - - -
SSIM↓ 0.78 0.92 0.22 0.60 - - -
F1-score↓ - - - - 0.77 0.19 0.74
0.5 L2mask↑ 0.01 0.01 0.23 0.17 - - -
SRmask↑ 0.00% 0.00% 99.96% 99.56% - - -
PSNR↓ 32.06 32.19 12.58 17.18 - - -
SSIM↓ 0.95 0.95 0.51 0.76 - - -
F1-score↓ - - - - 0.75 0.23 0.70
TABLE VIII: Compared with different feature processing strategies. The best result is marked in bold.
@@
Feature Enhancement HiSD 0.18 98.98% 16.75 0.78
AttGAN 0.13 96.06% 17.67 0.64
StarGAN 0.65 100.00% 7.17 0.24
AGGAN 0.29 100.00% 13.90 0.57
@@
The ability of TSDF to combine two defensive functions relies on the intensity separation mechanism, which is controlled by a threshold $\tau$. To find the optimal balance, we tested various threshold values, with the results presented in Table VII. The data show a clear functional conflict between the interruption and poisoning effects. At a low threshold such as 0.2, the framework dedicates more of the perturbation’s budget to interruption. This results in strong performance against deepfake models, achieving an L2mask of 0.67 on StarGAN, but leads to a weaker poisoning effect on face detectors with an F1-score of 0.87 on S3FD. Conversely, a high threshold like 0.5 allocates more space to poisoning, which significantly impairs face detectors by reducing the F1-score on S3FD to 0.70, but this is at the expense of weakening the interruption effect. Our experiments identify a threshold of 0.3 as the optimal balance, providing balanced performance in both interruption and poisoning across all models and detectors. This validates the effectiveness of the separation strategy.
To validate that the framework’s poisoning efficacy originates from our dedicated component, we evaluate an interruption-only variant of TSDF ($\tau$=0) in which the poisoning module is completely deactivated. The results are clear. The interruption-only perturbation of TSDF has almost no impact on the face detectors’ F1-score, which remained nearly the same as the undefended baseline. In contrast, the full TSDF framework significantly reduces the F1-score. This proves that the strong poisoning capability comes from the poisoning component and is not the effect of interruption.
@@
To isolate and evaluate the contribution of our Feature Enhancement strategy specifically to the interruption component, we conduct a dedicated ablation study. In this experiment, the poisoning module is deactivated across all test configurations to ensure that its effects do not interfere with the comparison of interruption performance. We then compare it with several alternative feature processing methods, using the Kullback-Leibler (KL) divergence and the Wasserstein distance [29]. As shown in Table VIII, the Feature Enhancement method consistently yields the best interruption performance. For example, for the StarGAN model, the strategy achieves an L2mask score of 0.65. These results are better than those of the KL divergence method, which results in an L2mask of 0.22. The baseline without enhancement performs even lower. This confirms that the Feature Enhancement strategy is more effective in creating powerful and disruptive perturbations. These results demonstrate the superior efficacy of our Feature Enhancement strategy, which produces a significantly more disruptive perturbation than the baselines. The high efficacy of this primary interruption is a key requirement for our two-stage paradigm, as it establishes the foundational defense that the poisoning component is designed to make persistent.

### v3

@@ IV-D Analysis of Defense Persistence
This section is dedicated to one of the contributions of this work, which is to validate the persistence of our framework. We conduct a direct comparative analysis that first establishes the failure of conventional SOTA defenses against an adaptive attack. We then demonstrate that our framework not only withstands the same attack but also enhances its defensive capabilities, thereby proving its superior persistence.
@@
The results presented in Table III reveal a critical vulnerability in conventional interruption-only methods. For instance, with the CMUA method, the average SSIM score on the CelebA dataset increases from 0.64 to 0.79 after retraining, while the average FID score falls from 209.46 to 127.42. The rising SSIM indicates that the output of perturbed examples has become more similar to the original images, while the falling FID confirms they have become more realistic and less distorted. Both metrics prove that the interruption effect has been largely nullified. This degradation provides clear quantitative evidence that adaptive attacks severely compromise conventional interruption-only defenses, proving their lack of persistence. This highlights the key vulnerability our TSDF method is designed to solve.
@@
To provide a direct comparison against this established baseline of failure, we subjected our TSDF method to the identical adaptive attack.
To validate the persistence of our framework, we conduct a direct comparative experiment against the SOTA interruption methods. To ensure a fair comparison, we evaluate all methods on the same two representative models, StarGAN and AttGAN. We simulate an adaptive attack by retraining these models on datasets protected by each respective defense. For TSDF, the retraining dataset is first processed by the poisoning mechanism. We then measure the interruption effectiveness of each defense before and after this adaptive attack to quantify their persistence.
The results presented in Table IV provide clear quantitative evidence of TSDF’s superior persistence. After retraining, the defenses of both CMUA and FOUND degrade significantly. CMUA’s SSIM score increase from 0.59 to 0.77, and its FID score drops from 243.00 to 122.46. These results confirm that the retrained models successfully learn to bypass these perturbations. In contrast, the output from the model retrained on TSDF-protected data become even more distorted. Its SSIM score decreases from 0.40 to 0.26, and its FID score increases from 285.56 to 344.12. This outcome is not because the interruption perturbation became stronger. It is direct evidence of the poisoning component’s success in degrading the quality of deepfake models.
This enhanced persistence is attributed to the efficacy of the poisoning component. As detailed in the model discrepancy analysis in Table V, retraining on TSDF data fundamentally destroys the deepfake model’s own generative capabilities. For instance, a StarGAN model’s FID score on clean inputs after this process increases from a baseline of 43.77 to 199.87. Since the model’s core ability to generate quality images is compromised, it cannot learn to properly adapt to or bypass the complex interruption perturbation.
@@ IV-E Visual and Qualitative Analysis
A crucial aspect of our framework is its ability to provide strong defense while remaining invisible to the human eye. Fig. 6 provides a direct visual comparison of an original image against images protected by CMUA, FOUND, and TSDF. The visual differences are nearly imperceptible, demonstrating the excellent concealment of our perturbation. This qualitative assessment is supported by the quantitative data in Table VI, which shows high PSNR and SSIM scores for the protected images across all datasets. For instance, on the CelebA dataset, TSDF achieves PSNR of 33.07 and SSIM of 0.89, demonstrating that our framework can provide a defense without compromising image imperceptibility.
@@ IV-F Ablation Study
We perform ablation studies to analyze the contribution of key components of the TSDF method. Specifically, we evaluate the effectiveness of the intensity separation mechanism and the Feature Enhancement strategy.
TABLE VII: Impact of Perturbation Threshold on Defense Performance. This table evaluates how different threshold settings affect the trade-off between the two defensive functions. The results show the performance across various deepfake models and face detectors. Blank entries indicate that the corresponding defensive effect is non-existent.
@@
0.00 L2mask↑ 0.18 0.20 0.38 0.18 - - -
SRmask↑ 99.59% 99.54% 100.00% 100.00% - - -
PSNR↓ 15.41 15.29 9.67 17.37 - - -
SSIM↓ 0.74 0.56 0.13 0.68 - - -
@@
0.02 L2mask↑ 0.18 0.21 0.40 0.19 - - -
SRmask↑ 99.45% 99.57% 100.00% 100.00% - - -
PSNR↓ 15.45 15.29 9.42 17.04 - - -
SSIM↓ 0.74 0.55 0.13 0.67 - - -
F1-score↓ - - - - 0.74 0.99 0.99
0.03 L2mask↑ 0.17 0.20 0.39 0.18 - - -
SRmask↑ 99.22% 99.56% 100.00% 100.00% - - -
PSNR↓ 15.83 15.52 9.47 17.47 - - -
SSIM↓ 0.75 0.56 0.13 0.68 - - -
F1-score↓ - - - - 0.75 0.84 0.99
0.04 L2mask↑ 0.12 0.17 0.37 0.18 - - -
SRmask↑ 99.15% 98.25% 100.00% 100.00% - - -
PSNR↓ 16.14 16.76 9.73 17.86 - - -
SSIM↓ 0.76 0.62 0.17 0.70 - - -
F1-score↓ - - - - 0.77 0.23 0.29
0.05 L2mask↑ 0.01 0.01 0.14 0.13 - - -
SRmask↑ 0.00% 0.00% 98.10% 99.22% - - -
PSNR↓ 30.69 31.50 14.80 22.10 - - -
SSIM↓ 0.94 0.94 0.46 0.83 - - -
F1-score↓ - - - - 0.97 0.03 0.05
TABLE VIII: Compared with different feature distance metrics strategies.
@@
Feature Enhancement HiSD 0.18 99.59% 15.41 0.74
AttGAN 0.20 99.54% 15.29 0.56
StarGAN 0.38 100.00% 9.67 0.13
AGGAN 0.18 100.00% 17.37 0.68
@@
The ability of TSDF to combine two defensive functions relies on the intensity separation mechanism, which is controlled by a threshold $\tau$. To find the optimal balance, we tested various threshold values, with the results presented in Table VII. The data show a clear functional conflict between the interruption and poisoning effects. At a low threshold such as 0.02, the framework dedicates more of the perturbation’s budget to interruption. This results in strong performance against deepfake models, achieving L2mask of 0.40 on StarGAN, but leads to a weaker poisoning effect on face detectors with F1-score of 0.99 on S3FD. Conversely, a high threshold like 0.05 allocates more space to poisoning, which significantly impairs face detectors by reducing the F1-score on S3FD to 0.03, but this is at the expense of weakening the interruption effect. Our experiments identify a threshold of 0.04 as the optimal balance, providing balanced performance in both interruption and poisoning across all models and detectors. This validates the effectiveness of the separation strategy.
To validate that the framework’s poisoning efficacy originates from our dedicated component, we evaluate an interruption-only variant of TSDF ($\tau$ = 0.00) in which the poisoning module is completely deactivated. The results are clear. The interruption-only perturbation of TSDF has almost no impact on the face detectors’ F1-score, which remained nearly the same as the undefended baseline. In contrast, the full TSDF method significantly reduces the F1-score. This proves that the strong poisoning capability comes from the poisoning component and is not the effect of interruption.
@@
To isolate and evaluate the contribution of our Feature Enhancement strategy specifically to the interruption component, we conduct a dedicated ablation study. In this experiment, the poisoning module is deactivated across all test configurations to ensure that its effects do not interfere with the comparison of interruption performance. We then compare it with several alternative feature distance metrics, using the Kullback-Leibler (KL) Divergence and the Wasserstein Distance [43]. As shown in Table VIII, the Feature Enhancement method consistently yields the best interruption performance. As shown in the case of StarGAN, the proposed strategy achieves an L2mask score of 0.38. This result surpasses that of the KL Divergence method, where the corresponding L2mask value is only 0.22. The baseline without enhancement performs even lower. This confirms that the Feature Enhancement strategy is more effective in creating powerful and disruptive perturbations. These results demonstrate the superior efficacy of our Feature Enhancement strategy, which produces a significantly more disruptive perturbation than the baselines. The high efficacy of this primary interruption is a key requirement for our two-stage paradigm, as it establishes the foundational defense that the poisoning component is designed to make persistent.